# A Deep Learning Framework for Disentangling Triangle Singularity and Pole-Based Enhancements

### v2

@@ Acknowledgements.
DAOC acknowledges the support of the DOST-SEI Undergraduate Merit Scholarship Program. VAAC acknowledges the support of DOST-ASTHRDP. The authors acknowledge the helpful insights of Leonarc Michelle Santos and Mao-Jun Yan.
@@ Appendix A Sample Datasets
On the other hand, Figure 12 shows the three pole configurations considered in the study. The P01 and P02 training sets are the label used for line shapes with an isolated pole in the $[bt]$-sheet and $[tb]$-sheet, respectively. The general features of these line shapes, specifically the appearance of peak relative to the second threshold ($\Sigma_{c}^{+}\bar{D}^{0}$), is shared by the datasets. Notice that the peak is relatively weak in P02 compared to the other classifications, which is also demonstrated in Figure 4. The suppressed strength of an isolated pole in the $[tb]$-sheet is due to the fact that it is located in the farthest Riemann sheet among the unphysical sheets considered. Furthermore, the peak is also restricted by the unitarity of the $S$-matrix between the two thresholds, suppressing the signal even further. Lastly, the P03 dataset has the most non-trivial line shapes among the different interpretations considered in this work. Recall that two poles are used in P03, one in $[bt]$ and the other one in $[bb]$. In some cases, these poles can produce two peaks in the line shapes and is easily distinguishable from the other classifications. However, with this configuration, there are cases when the $[bt]$ and the $[bb]$ poles are both below the second threshold. In this situation, only one peak is observed associated with the $[bt]$ pole while the $[bb]$ pole serves to lessen the cusp around the second threshold. In fact, when these poles have exactly the same position, the peak will have a symmetric shape despite the presence of the second threshold. This last pole-based classification is expected to produce line shapes that is quite similar with the triangle shapes. This observation makes the line shape classification a non-trivial task without the use of an automated pattern recognition method such as machine learning.

### v3

@@ Acknowledgements.
DAOC acknowledges the support of the DOST-SEI Undergraduate Merit Scholarship Program. VAAC acknowledges the support of DOST-ASTHRDP. The authors acknowledge the helpful suggestions of Leonarc Michelle Santos and Mao-Jun Yan. DLBS acknowledges the insightful discussions with Prof. Tetsuo Hyodo and Prof. Atsushi Hosaka during the long-term workshop on HHIQCD2024 at the Yukawa Institute for Theoretical Physics (YITP-T-24-02).
@@ Appendix A Sample Datasets
On the other hand, Fig. 10 shows the three pole configurations considered in the study. The P01 and P02 training sets are the label used for line shapes with an isolated pole in the $[bt]$-sheet and $[tb]$-sheet, respectively. The general features of these line shapes, specifically the appearance of peak relative to the second threshold ($\Sigma_{c}^{+}\bar{D}^{0}$), is shared by the datasets. Notice that the peak is relatively weak in P02 compared to the other classifications, which is also demonstrated in Fig. 4. The suppressed strength of an isolated pole in the $[tb]$-sheet is due to the fact that it is located in the farthest Riemann sheet among the unphysical sheets considered. Furthermore, the peak is also restricted by the unitarity of the $S$-matrix between the two thresholds, suppressing the signal even further. Lastly, the P03 dataset has the most non-trivial line shapes among the different interpretations considered in this work. Recall that two poles are used in P03, one in $[bt]$ and the other one in $[bb]$. In some cases, these poles can produce two peaks in the line shapes and is easily distinguishable from the other classifications. However, with this configuration, there are cases when the $[bt]$ and the $[bb]$ poles are both below the second threshold. In this situation, only one peak is observed associated with the $[bt]$ pole while the $[bb]$ pole serves to lessen the cusp around the second threshold. In fact, when these poles have exactly the same position, the peak will have a symmetric shape despite the presence of the second threshold. This last pole-based classification is expected to produce line shapes that is quite similar with the triangle shapes. This observation makes the line shape classification a non-trivial task without the use of an automated pattern recognition method such as machine learning.
Appendix B Uncertainty Estimation
In this appendix, we briefly describe how the snapshot ensemble is used to estimate the uncertainty of our trained DNN. During training, we saved the state of our DNN models from epoch 950 to epoch 1050, creating a snapshot ensemble of 100 different DNN instances. This ensemble is utilized to estimate the uncertainty of the inference results given the experimental data. Our approach is schematically outlined in Fig. 11. We first generated 3000 inference samples by bootstrapping the given experimental data. These 3000 samples were obtained using a uniform distribution, with extreme values marked by the error bars and energy uncertainty present in the experimental data. These samples were then fed directly into the 100 snapshot states of the DNN, and the number of classification outputs was counted. From the distribution of the DNN’s outputs, we can estimate the mean and standard deviations for each classification. This method provides a comprehensive measure of uncertainty, which is crucial for understanding the reliability of the model’s predictions and for making informed decisions based on the inference results.
Figure 11: Schematic diagram of the snapshot ensemble method to estimate the uncertainty of the DNN’s output.